Find Y* and C*

Y= C+ I+G

C=25+6Y*1/2

I=16

G= 14


1
Expert's answer
2021-04-19T18:46:32-0400
"solution"

"Y = C + I + G \\
C = 25 + 6Y
^\frac{1}{2}
\\
I = 16 \\
G = 14"


"Substituting\ C, I, G\ values :\\Y = 25 + 6Y^\frac{1}{2} + 16 + 14\\Y = 6Y^\frac{1}{2} + 55\\ Y - 55 = 6Y^\frac{1}{2 }\\"

"But\ we \ know\ that :\\Y = (Y
^\frac{1}{2}
)
2
.\\ Let \ Y
^\frac{1}{2} = a, \ and \ we\ can\ rewrite\ above \ in \ the\ following \ way:\\
a
^2 = 55 + 6a \\ a^
2 − 6a =55"


"using \ the\ quadratic\ formula\ ,we \ should\ be\ able\ to\ find \ that:\\a \ is\ either \ a= 11 \ or\ 
a= −5. \\This\ allows\ us\ to \ find\ that:\\ Y = 121. \\Note\ that\ a = −5 \ doesn’t \ make\ economic\ sense.\\
Given \ that\ Y
 = 121, we\ can\ go\ back , \ and\ find \ that\\ C =25+66\\C= 91"